A man spent one fifth of his money at the butchers and one third of his money at the supermarket. He had £5.60 left. How much did he start with?


1
Expert's answer
2021-05-21T03:30:25-0400

"\frac{1}{5}" spent at the butchers


"\frac{1}{3}" spent at the supermarket


Calculate what fraction of the money was left:


"1 - ( \frac{1}{5} + \frac{1}{3} ) =" "\frac{7}{15}"


"\frac{7}{15}" "x =" 5.60


"x =" 12


Therefore the starting amount was 12 pounds